40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 25 19:42:12 2014, ericq, Update, General, Vent progress 
    Reply  Fri Sep 26 12:25:34 2014, ericq, Update, General, Vent progress 
       Reply  Fri Sep 26 16:10:14 2014, ericq, Update, General, Vent update 
          Reply  Fri Sep 26 17:13:39 2014, ericq, Update, General, Vent update 
             Reply  Mon Sep 29 12:47:51 2014, ericq, Update, General, Vent update 
                Reply  Mon Sep 29 17:10:51 2014, ericq, Update, General, Vent update 
                   Reply  Mon Sep 29 18:12:24 2014, ericq, Update, General, Vent update 
                      Reply  Tue Sep 30 11:53:29 2014, ericq, Update, General, Vent update 
                         Reply  Tue Sep 30 16:18:57 2014, Steve, Update, VAC, pump down #78 has started pd78hastarted.png
                            Reply  Tue Sep 30 18:02:53 2014, Steve, Update, VAC, pump down #78 stops at 320 Torr 3h.png
                               Reply  Wed Oct 1 10:21:19 2014, Steve, Update, VAC, pump down #78 stops at 3.5 Torr pd78secondStop.png
                                  Reply  Wed Oct 1 16:26:53 2014, Steve, Update, VAC, pump down #78 completed pd78completedin25hrs.png
                                     Reply  Thu Oct 2 09:23:23 2014, Steve, Update, VAC, vent 77 vent77SRMmoved.png
                                        Reply  Fri Oct 3 10:10:37 2014, Steve, Update, VAC, cold cathode gauge reading switched CC1switched.png
                                           Reply  Fri May 8 14:30:09 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now cc1bake95C.jpg
                                              Reply  Tue May 12 14:57:52 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now 
Message ID: 10541     Entry time: Thu Sep 25 19:42:12 2014     Reply to this: 10544
Author: ericq 
Type: Update 
Category: General 
Subject: Vent progress 

[q, Jenne, Manasa]

ELOG Outline

  • Aligned arms
  • Took a bunch of photos of ITMY chamber. 
  • Used Crystallaser to reverse trace POY beam path
  • Realized real POY flashes were visible
  • Tried adjusted POY steering to give us enough room to move SRM forward 3 in
  • Not enough steering room
  • Placed beam centering targets just before OM3 and after OM4, on far side of BS table
  • Twisted SR2 by ~5 degrees
  • Moved SRM laterally ~3 inches (beam had been hitting optic center, then suspension cage after SR2 twist)
  • Moved SRM foward about 3 inches, adjusted angle for coarse retroreflection
  • Measured distances between suspensions
  • Tried running Gabriele's distance reconstruction code, results not looking so good; redundancy checks are off by ~1cm
  • Started roughly repositioning OM1 and OM2 to get through beam targets
  • Closed light doors
  • ITMY pointing has gone bad, no green or IR resonance to be seen, still have IR and Green in Xarm, so TTs, BS are ok
ELOG V3.1.3-